OpenStudy (chihiroasleaf):

Find the value of a and b so that \[x ^{4}-ax ^{3}-\left( 6a+5b \right)x ^{2}+abx+144\] can be divided by \[x ^{2}+6x+8\]! I've tried to use long division but it seems doesn't work...,

hartnn (hartnn):

hint : if that expression can be divided by that quadratic equation, then roots of that quadratic equation satisfy that 4th degree expression.

hartnn (hartnn):

say if you get 2 roots as p and q , then put x=p, and x=q in that 4th degree equation, to get 2 equations in a and b which can be solved simultaneously ....

OpenStudy (chihiroasleaf):

the roots of \[x ^{2}+6x+8\] are -2 and -4 what is f(-2) and f(-4) ? are f(-2) and f(-4) equals zero?

OpenStudy (chihiroasleaf):

I mean the roots of quadratic equation \[x ^{2}+6x+8=0\]

OpenStudy (anonymous):

x^2 +6x+8 =(x+2) (x+4) So, Let f(x)= x^4 -ax^3 -(6a+5b)x^2+abx+144

hartnn (hartnn):

yes, because f(x) can be divided without remainder , so, f(-2)=0 , f(-4)=0.

OpenStudy (anonymous):

Since f(x) is divided by x^2+6x+8 f(-2)=0 f(-4)=0 U will get 2 equations and u need to solve them

Parth (parthkohli):

The remainder theorem says that the remainder of \(\dfrac{p(x)}{q(x)}\) is \(p(a)\) where \(a\) is the root of \(q(x)\).

OpenStudy (chihiroasleaf):

I just know the remainder theorem for division by linear expression like (x-a) so it also hold for quadratic expression?

hartnn (hartnn):

remember that quadratic expression is product of 2 linear expressions, so yes.

OpenStudy (chihiroasleaf):

hhmm..., if (x-a)(x-b) divide f(x) ..., is (x-a) divide f(x) and (x-b) divide f(x) ? if the remainder theorem holds, then it should be true, right? I'm confused with this.., so I'm not sure in applying remainder theorem before...

hartnn (hartnn):

example, if (x-4)(x+4) divide f(x) [let f(x)=x^2-16] then isn't it true that both x-4 and x+4 divide x^2-16 and f(4) = f(-4) =0.

OpenStudy (chihiroasleaf):

I see.... :D then, I get this.. f(-2) = 10a + 5b -ab = -80 f(-4) = 22a+5b+ab = -100 I add both equation, then I get 16a+5b = -90... and I get stuck... :(

OpenStudy (chihiroasleaf):

I think I've an idea to solve it..., thanks a lot.... :)

hartnn (hartnn):

u sure those equations are correct ? http://www.wolframalpha.com/input/?i=solve+10a+%2B+5b+-ab+%3D+-80+%2C+22a%2B5b%2Bab+%3D+-100

OpenStudy (chihiroasleaf):

ahh.., I made mistake in calculation it should be \[f(-2) = 10a+5b-ab=-80\] \[f(-4)= 22a+5b-ab=-100\]

hartnn (hartnn):

then (-5/3 , -19/2)

OpenStudy (chihiroasleaf):

yess..., I get it..., thanks a lot... :D

hartnn (hartnn):

welcome ^_^